# reference.bib

### v1

@@ IV-A1 Dataset
The primary dataset used in our experiments is the MetaFruit dataset [li2025metafruit], a multi-class fruit detection benchmark consisting of 4,248 high-resolution images. It includes 248,015 manually annotated instances, collected from diverse orchards in Michigan and California during the 2022 and 2023 growing seasons. The dataset used in this experiment covers three fruit categories: apples, oranges, lemons, with each image annotated using precise bounding boxes and class labels. MetaFruit poses significant challenges due to factors such as varying lighting conditions, occlusions, and clustered fruit appearances, making it well-suited for evaluating the proposed VLLFL framework in realistic agricultural scenarios. Notably, MetaFruit is the largest publicly available multi-class fruit dataset to date, surpassing prior datasets by more than an order of magnitude in the number of labeled instances.
To further evaluate the generalization capability of VLLFL, we also employ the Harmful Animal in the Field dataset [harmfulanimalinthefield_dataset]. This dataset contains 9,849 images aimed at detecting animals that pose potential threats to agricultural environments. It comprises 31 distinct classes, including buffalo, chicken, deer, elephant, goat, monkey, pig, and others. In our experiments, we randomly select 4 classes for evaluation to simulate varied detection tasks.
@@ IV-A3 Baseline method
In our evaluation, we adopt Federated Context Optimization (FedCoOp) [guo2023promptfl] as the baseline method to highlight the advantages of the proposed VLLFL framework. FedCoOp learns continuous prompt embeddings to enhance the performance of vision-language models (VLMs) across federated tasks. Once optimized, these prompt embeddings remain fixed during inference. This is a key difference between FedCoOp and our proposed approach. In contrast, VLLFL is designed to train a context-aware prompt generator capable of dynamically producing optimized prompts for different classes, rather than relying on static, pre-trained embeddings. This flexibility allows VLLFL to adapt to diverse task distributions across clients and achieve improved generalization in federated settings.

### v2

@@ IV-A1 Dataset
The primary dataset used in our experiments is the MetaFruit dataset [45], a multi-class fruit detection benchmark consisting of 4,248 high-resolution images. It includes 248,015 manually annotated instances, collected from diverse orchards in Michigan and California during the 2022 and 2023 growing seasons. The dataset used in this experiment covers three fruit categories: apples, oranges, lemons, with each image annotated using precise bounding boxes and class labels. MetaFruit poses significant challenges due to factors such as varying lighting conditions, occlusions, and clustered fruit appearances, making it well-suited for evaluating the proposed VLLFL framework in realistic agricultural scenarios. Notably, MetaFruit is the largest publicly available multi-class fruit dataset to date, surpassing prior datasets by more than an order of magnitude in the number of labeled instances.
To further evaluate the generalization capability of VLLFL, we also employ the Harmful Animal in the Field dataset [46]. This dataset contains 9,849 images aimed at detecting animals that pose potential threats to agricultural environments. It comprises 31 distinct classes, including buffalo, chicken, deer, elephant, goat, monkey, pig, and others. In our experiments, we randomly select 4 classes for evaluation to simulate varied detection tasks.
@@ IV-A3 Baseline method
In our evaluation, we adopt Federated Context Optimization (FedCoOp) [47] as the baseline method to highlight the advantages of the proposed VLLFL framework. FedCoOp learns continuous prompt embeddings to enhance the performance of vision-language models (VLMs) across federated tasks. Once optimized, these prompt embeddings remain fixed during inference. This is a key difference between FedCoOp and our proposed approach. In contrast, VLLFL is designed to train a context-aware prompt generator capable of dynamically producing optimized prompts for different classes, rather than relying on static, pre-trained embeddings. This flexibility allows VLLFL to adapt to diverse task distributions across clients and achieve improved generalization in federated settings.